OpenStudy (anonymous):

Let f(x) = x2 – 16. Find f–1(x).

OpenStudy (anonymous):

assuming that x2= x^2 and (f-1)(x) f(x) = x^2 -16 (f-1)(x) = x^2-x-16

OpenStudy (amistre64):

to find an inverse, a good trick is to swap out x and y and resolve for y

OpenStudy (amistre64):

y = x^2 - 16 ; swap out x and y x = y^2 - 16 resolve for y